in a triangle with the side lengths 5, 7, 9 length units, the largest angle x is

A) x>90

B) x=90

C) x<90


1
Expert's answer
2021-05-13T06:23:17-0400

We have given the sides of triangle "5,7,9."


Largest angle would be on the opposite of largest side,


We know,


"cosA = \dfrac{b^2+c^2-a^2}{2bc}"


"cosA = \dfrac{5^2+7^2-9^2}{2 \times 5 \times 7}"


"cosA = -\dfrac{21}{70}"


"A = 107.45^{\circ}"


Hence the angle will be greater than "90^{\circ}"